# Optimal network sizes for most robust Turing patterns

### v1

@@ Results
Figure 6: Relationships between network size $N$ and radius $\gamma=\sigma\sqrt{N}$. $10^{3}$ random matrices are sampled for each $N$ and $\gamma$ pair. (A) Percentage of stable random matrices without diffusion. (B) Percentage of previously stable random matrices turning unstable with diffusion. ’Hopf’ indicates Turing-Hopf. (C) Percentage of random matrices leading to Turing instabilities (Turing I, II, and Turing-Hopf). (D) Percentage of random matrices leading to Turing I instabilities.
@@
Without having an experimental understanding of the variability of interaction strength between nodes, an optimal network size should produce Turing I instabilities for a wide range of variances. Such a network size would be most robust to changes in interaction strength, which is a desirable outcome from a biological perspective. To find the exact network size that is the most optimal, we plotted the occurrence of Turing instabilities for different network sizes and variances from $N=3$ to $100$ (Fig. 7A). Based on the heat map, the majority of the Turing instabilities arose just above $\gamma=1$, as we should expect since $\gamma>\!\!>1$ would likely result in an unstable system without diffusion, but $\gamma<\!\!<1$ would make it harder for a stable system to become unstable with diffusion. In fact, the occurrence of Turing is ”pinned” to radius $\gamma=1$ for large $N$, and the percentage density then follows closely the circular law $\sigma^{2}\sim 1/N$.
@@
Is there an optimal network size for highest robustness of the occurance of Turing patterns? We expect such an $N_{\text{opt}}$ as there are no Turing patterns for $N=2$ and for $N\rightarrow\infty$. Turing patterns disappear as the eigenvalue distribution becomes a step function, allowing no instability with diffusion. To confirm our intuition, we then calculated the percentage of Turing I for each network size from the heat map (Fig. 7B), demonstrating that $N_{\text{opt}}=5$ is the most optimal network size based on its highest percentage of Turing patterns. Thus, for a network with two diffusible species, having additional three immobile nodes gives the optimal network size for Turing instabilities. Compared to $N=3$ with one non-diffusible species (4.73%), Turing instabilities are three times more likely to occur for $N=5$ (13.86%).
@@
Figure 7: Optimal network size for highest Turing robustness. (A) Heat map of percentage occurrence of Turing I in random matrices of different networks sizes $N$ and variances $\sigma^{2}$ for $\gamma=1$ and zero off-diagonal element sparsity (or connectivity $C=1$). (B) Corresponding percentage shares of each network size $N$. For our parameters, the optimal network size is $N=5$. The horizontal axis is set to a linear scale (logarithmic scale in inset).
@@
While our random matrices are fully connected, it was suggested that biological networks tend to be sparsely connected due to its evolutionary advantage in preserving robustness [31]. Let us assume that the probability of observing $k$ connections is binomially distributed $\binom{n}{k}p^{k}(1-p)^{n-k}$ where $n$ the number of matrix elements to consider (’trials’) such as $n=N(N-1)$ for the number of off-diagonal matrix elements, and $p$ the probability for having a connection (’success’). As the variance $np(1-p)$ is largest for $p=0.5$, corresponding to the maximum entropy distribution, we would expect on average 50% connections. This corresponds to the maximum number of possible network structures found by evolution. However, does sparcity also translate into higher Turing robustness? We incorporated sparsity but found that its effect is minor, without qualitatively changing our results (see Figs. S2 and S3 in Supplementary Materials). The peak value of the Turing I percentage slightly reduces with sparcity ($\approx 14$, $12$, and $10$% for respective sparcities 00, $25$, and $50$%), while the corresponding optimal network sizes increase slightly ($N_{\text{opt}}=5$, $7$, and $8$ for above sparcities). Hence, overall the robustness of Turing is reduced with sparcity, pointing to feedback being important for pattern formation. Note the circular law also applies to sparcity, but needs modification. The radius is just rescaled to $\gamma=\sigma\sqrt{NC}$ where connectivity $C=p$ is the probability for having an off-diagonal matrix element [25]. This points to our results being broadly representative for a large range of random Turing networks.
@@
A major constraint on Turing patterns is that the inhibitor has to diffuse much faster than the activator. How does having immobile nodes change this constraint? Previous work on small networks hinted towards a softening of this constraint [23, 22, 26]. By varying the diffusion of the two diffusible species modeled in our random reaction-diffusion systems, we sampled systems of $N=2$, $3$, $5$, and $50$ nodes, with $N=2$ not supporting Turing for our random matrices with equal degradation of $-1$ for all nodes (see proof in Methods). Importantly, we found that differential diffusivity is important for Turing I instabilities to arise for very few immobile nodes, but not for systems with many immobile nodes: Based on Fig. 8, a network of $N=3$ with one immobile node has nearly zero percent Turing instabilities arising at equal diffusivity (i.e. at the identity line $D_{1}/D_{2}=1$). As the diffusion ratio deviates from $1$, Turing instabilities become more likely for $N=3$, creating a distinct gap around the identity line. Notably, the plot is symmetric with respect to the identity line for all $N$ since the assignment to activator or inhibitor is arbitrary. In contrast, in the presence of more immobile nodes, Turing instabilities can arise at equal diffusivity: As network size increases, the gap at the identity line becomes more narrow, indicating Turing instabilities occurring at a wider combination of the diffusion parameters (see also Fig. S4 in the Supplementary Materials). For the a large network size of $N=50$, we observed a similar occurrence of Turing instability for the majority of the diffusion parameter combinations, except for when the diffusion parameters are both extremely low, approximating the case of no diffusion. Since the Turing instability is a diffusion-driven instability, it is reasonable to have no Turing instability arising here. Additionally, there is no ’dark line’ (indicating zero or low percentage) at the very bottom or very left of all surface plots, indicating that a Turing instability can arise in systems with only one diffusible species. The absence of differential diffusion is even stronger when consider all Turing cases (Turing I, II, and Turing-Hopf) - see Fig. S5 of the Supplementary Materials. In summary, having numerous immobile nodes significantly relaxes the constraints from Turing conditions.
@@
Figure 8: Effect of diffusion constants on Turing pattern formation. Heat map of percentage occurrence of Turing I in random matrices for different diffusion parameters, $D_{1}$ and $D_{2}$: (A) $N=2$, $\sigma^{2}=0.5$; (B) $N=3$, $\sigma^{2}=0.33$; (C) $N=5$, $\sigma^{2}=0.2$ (D) $N=50$, $\sigma^{2}=0.02$. For increasing $N$, constraints on diffusion constants being different vanish.
@@ Discussion
Applying Robert May’s circular law [24] to cases without and with diffusion of two designated species, in line with Turing’s original model [4], we identified an optimal network size for maximal robustness, reaching 13.86% of sampled matrices. This value is orders of magnitude higher than previous estimates for small networks [22]. Our results are robust even when sparcity is introduced, leading only to minor changes in the observed Turing percentage or optimal network size. Importantly, an optimum in Turing robustness emerges in our model due to the tradeoff between maximizing stability without diffusion, which favors small networks centered around eigenvalues on the negative real axis, and maximizing instability with diffusion, which favors large networks for eigenvalues to reach positive real values. According to the circular law, small networks have small radii for the distribution of their eigenvalues, while large networks have large radii. For increasingly large Turing networks, we further found that differential diffusion constants are not required, in line with studies on Turing topologies [23]. Having many immobile nodes effectively allows for delays and hence different diffusion constants. We hope these results shed new light on Turing mechanisms in developmental systems and help make Turing mechanisms more relevant to quantitative biology.
@@
To make our model interpretable, we used a number of simplifying assumptions, resulting in limitations of our predictions. A major simplification is our diagonal matrix elements, which are set to a constant value for all species, representing equal degradation rates for all species. This led to the convenient behavior that all eigenvalues are centered around a single point on the negative real axis, following Robert May’s circular law. Without that assumption, the eigenvalue distributions would still be restricted by circles, but now following the less restrictive Gerschgorin theorem. Based on the latter, the radius scales as $\sim N$, and not $\sim\sqrt{N}$ as in May’s circular law. This leads not only to larger circles but also to unions of overlapping circles, restricting the real parts less efficiently [33]. Using the same degradation rates also does not allow for flexibility in the levels of self-activation and inhibition often observed in biological pathways. Choosing degradation rates from a distribution introduces correlations among corresponding off-diagonal matrix elements $J_{ij}$ and $J_{ji}$ and hence different shapes of eigenvalue distributions [25].
@@ Methods
For Fig. 3 we sampled $10^{7}$ parameter combinations for each network topology using Latin hypercube sampling lhs function from pyDOE library, using a loguniform distribution with the following bounds: $V_{i}:\ (0.1,10)$, $K_{ij}:\ (0.01,1)$, $b_{i}:\ (0.001,0.1)$, $\mu_{i}:\ (0.01,1)$ (see Eq. 2). The diffusion constants for Species 1 and 2 were fixed to $1$ and $10$, and the Hill coefficient was set to $n_{ij}=2$ throughout. For each parameter set we calculated the unique steady states using $10$ different initial guess using the Newton-Raphson method (function root with ’hybrid’ setting). In Figs. 6 and 7 we used $10^{3}$ random matrices for each $N$ and $\gamma$ pair, and in Fig. 8 we used $10^{3}$ random matrices for each $D_{1}$ and $D_{2}$ pair.
@@
Fitting a beta distribution to the data of the Jacobian matrix elements in Fig. 3 involves estimating the distribution’s parameters, typically using the method of moments or maximum likelihood estimation (MLE). The method of moments involves calculating the sample mean and variance to derive initial estimates for the shape parameters $\alpha$ and $\beta$. MLE, on the other hand, maximizes the log-likelihood function, which measures how well the beta distribution with specific parameters explains the observed data, to find the best-fitting parameters. Both methods aim to provide a beta distribution that accurately represents the underlying data distribution. Here, we used the method of moments to obtain an initial guess for the parameter values of $\alpha$ and $\beta$, which we then used as an initial guess for the maximisation of the log-likelihood function.
@@
Hence, the last of Turing’s conditions cannot be fulfilled for two nodes, which is relevant for Figs. 7A and 8A.

### v2

@@ Results
Figure 6: Relationships between network size $N$ and radius $\gamma=\sigma\sqrt{N}$. $10^{3}$ random matrices are sampled for each $N$ and $\gamma$ pair. (A) Percentage of stable random matrices without diffusion. (B) Percentage of previously stable random matrices turning unstable with diffusion. ’Hopf’ indicates Turing-Hopf. (C) Percentage of random matrices leading to Turing instabilities (Turing I, II, and Turing-Hopf). (D) Percentage of random matrices leading to Turing I instabilities. The key point is $\gamma=1$, while for larger $\gamma$ values the statistics become unreliable.
@@
Without having an experimental understanding of the variability of interaction strength between nodes, an optimal network size should produce Turing I instabilities for a wide range of variances. Such a network size would be most robust to changes in interaction strength, which is a desirable outcome from a biological perspective. To find the exact network size that is the most optimal, we plotted the occurrence of Turing instabilities for different network sizes and variances from $N=2$ to $100$ (Fig. 7A). Based on the heat map, the majority of the Turing instabilities arose just above $\gamma=1$, as we should expect since $\gamma>\!\!>1$ would likely result in an unstable system without diffusion, but $\gamma<\!\!<1$ would make it harder for a stable system to become unstable with diffusion. In fact, the occurrence of Turing is ”pinned” to radius $\gamma=1$ for large $N$, and the percentage density then follows closely the circular law $\sigma^{2}\sim 1/N$.
@@
Is there an optimal network size for highest robustness of the occurance of Turing patterns? We expect such an $N_{\text{opt}}$ as there are no Turing patterns for $N=2$ and for $N\rightarrow\infty$. Turing patterns disappear as the eigenvalue distribution becomes a step function, allowing no instability with diffusion. To confirm our intuition, we then calculated the percentage of Turing I for each network size from the heat map (Fig. 7B), demonstrating that $N_{\text{opt}}=5$ is the most optimal network size based on its highest percentage of Turing patterns. Thus, for a network with two diffusible species, having additional three immobile nodes gives the optimal network size for Turing instabilities. Compared to $N=3$ with one non-diffusible species (4.73%), Turing instabilities are three times more likely to occur for $N+\text{opt}=5$ (13.86%).
The peak is however not at the first allowed $N$ value, i.e. $3$, but at a higher value, pointing towards another principle for this maximum in Turing frequency. For finite-$N$ random matrices and $\gamma\gtrsim 1$, the scaling of the eigenvalue density, projected onto the real axis, near $x=0$ can be approximated as
$\rho(0)\sim\frac{1}{\pi}\sqrt{\gamma^{2}-1}\cdot\left(1-\frac{1}{2\gamma^{2}N}+\mathcal{O}\left(\frac{1}{N^{2}}\right)\right)$
(see the Supplementary Text for details). For small $N$, the density of marginal stable matrices i.e. near $x=0$, is suppressed. This is due to eigenvalues being correlated and biased toward the center of the eigenvalue circle due to finite $N$ effects. As a result, there are less Turing cases for small $N$ than expected based on the asymptotic form of Robert May’s circular law. This ultimately leads to the value $N_{\text{opt}}=5$.
@@
Figure 7: Optimal network size for highest Turing robustness. (A) Heat map of percentage occurrence of Turing I in random matrices of different network sizes $N$ and variances $\sigma^{2}$ for $\gamma=1$ and zero off-diagonal element sparsity (or connectivity $C=1$). (B) Corresponding percentage shares of each network size $N$. For our parameters, the optimal network size is $N=5$. The inset provides a more detailed view of network sizes from 2 to 8 for finer resolution.
@@
While our random matrices are fully connected, it was suggested that biological networks tend to be sparsely connected due to its evolutionary advantage in preserving robustness [36]. Let us assume that the probability of observing $k$ connections is binomially distributed $\binom{n}{k}p^{k}(1-p)^{n-k}$ where $n$ the number of matrix elements to consider (’trials’) such as $n=N(N-1)$ for the number of off-diagonal matrix elements, and $p$ the probability for having a connection (’success’). As the variance $np(1-p)$ is largest for $p=0.5$, corresponding to the maximum entropy distribution, we would expect on average 50% connections. This corresponds to the maximum number of possible network structures found by evolution. However, does sparsity also translate into higher Turing robustness? We incorporated sparsity but found that its effect is minor, without qualitatively changing our results (see Figs. S2 and S3 in the Supplementary Figures). The peak value of the Turing I percentage slightly reduces with sparsity ($\approx 14$, $12$, and $10$% for respective sparsities 00, $25$, and $50$%), while the corresponding optimal network sizes increase slightly ($N_{\text{opt}}=5$, $7$, and $8$ for above sparsities). Hence, overall the robustness of Turing is reduced with sparsity, pointing to feedback being important for pattern formation. Note the circular law also applies to sparsity, but needs modification. The radius is just rescaled to $\gamma=\sigma\sqrt{NC}$ where connectivity $C=p$ is the probability for having an off-diagonal matrix element [30]. This points to our results being broadly representative for a large range of random Turing networks.
Additionally, we tested full Gaussian matrices with Gaussian-distributed diagonal elements (allowing self-interactions), both with and without self-degradation ($-1$), as well as anti-symmetric matrices with diagonal elements fixed at -1, as in the original setup. The anti-symmetry is designed to capture correlations arising from potential mass-conserved reactions. The full Gaussian matrices with fixed degradation produce an optimal network size (see Fig. S4-S5), whereas the anti-symmetric matrices with $-1$ on the diagonals do not produce Turing patterns at any network size or variance, in line with known theorems [37]. This is because the eigenvalues of real anti-symmetric matrices are purely imaginary or zero, and adding -1 to the diagonal shifts them along the real axis, leaving no positive real parts to enable Turing instabilities. For the full Gaussian matrices, Turing patterns are possible for $N=2$ when self-interactions are present, as degradation is no longer fixed at $-1$ for all nodes. However, when there is no fixed degradation of $-1$, no clear peak in network size emerges, as the eigenvalue distribution becomes more diffuse without a dominant eigenvalue near $-1$ to stabilize the system.
@@
A major constraint on Turing patterns is that the inhibitor has to diffuse much faster than the activator. How does having immobile nodes change this constraint? Previous work on small networks hinted towards a softening of this constraint [28, 23, 31]. By varying the diffusion of the two diffusible species modeled in our random reaction-diffusion systems, we sampled systems of $N=2$, $3$, $5$, and $50$ nodes, with $N=2$ again not supporting Turing for our random matrices with equal degradation of $-1$. Importantly, we found that differential diffusivity is important for Turing I instabilities to arise for very few immobile nodes, but not for systems with many immobile nodes: Based on Fig. 8, a network of $N=3$ with one immobile node has nearly zero percent Turing instabilities arising at equal diffusivity (i.e. at the identity line $D_{1}/D_{2}=1$). As the diffusion ratio deviates from $1$, Turing instabilities become more likely for $N=3$, creating a distinct gap around the identity line. Notably, the plot is symmetric with respect to the identity line for all $N$ since the assignment to activator or inhibitor is arbitrary. In contrast, in the presence of more immobile nodes, Turing instabilities can arise at equal diffusivities: As network size increases, the gap at the identity line becomes more narrow, indicating Turing instabilities occurring at a wider combination of the diffusion parameters (see also Fig. S6 in the Supplementary Figures). For the large network size of $N=50$, we observed a similar occurrence of Turing instability for the majority of the diffusion parameter combinations, except for when the diffusion parameters are both extremely low, approximating the case of no diffusion. Since the Turing instability is a diffusion-driven instability, it is reasonable to have no Turing instability arising here. Additionally, there is no ’dark line’ (indicating zero or low percentage) at the very bottom or very left of all surface plots, indicating that a Turing instability can arise in systems with only one diffusible species. The absence of differential diffusion is even stronger when consider all Turing cases (Turing I, II, and Turing-Hopf) - see Fig. S7 of the Supplementary Figures. In summary, having numerous immobile nodes significantly relaxes the constraint of differential diffusivity from the classic Turing conditions.
@@
Figure 8: Effect of diffusion constants on Turing pattern formation. Heat map of percentage occurrence of Turing I in random matrices for different diffusion parameters, $D_{1}$ and $D_{2}$: (A) $N=2$, $\sigma^{2}=0.5$; (B) $N=3$, $\sigma^{2}=0.33$; (C) $N=5$, $\sigma^{2}=0.2$ (D) $N=50$, $\sigma^{2}=0.02$. For increasing $N$, the constraint of differential diffusivity begins to vanish.
@@ Discussion
Applying Robert May’s circular law [29] to cases without and with diffusion of two designated species, in line with Turing’s original model [4], we identified an optimal network size for maximal robustness, reaching 13.86% of sampled matrices. This value is orders of magnitude higher than previous estimates for small networks [23]. Our results are robust even when sparsity is introduced, leading only to minor changes in the observed Turing percentage or optimal network size. Importantly, an optimum in Turing robustness emerges in our model due to the tradeoff between maximizing stability without diffusion, which favors small networks with eigenvalues having negative real parts, and maximizing instability with diffusion, which favors large networks with eigenvalues having positive real parts. According to the circular law, small networks have small radii for the distribution of their eigenvalues, while large networks have large radii. For increasingly large Turing networks, we further found that differential diffusivity is not required, in line with studies on Turing topologies [28]. Having many immobile nodes effectively allows for delays and hence different diffusion constants. We hope these results shed new light on Turing mechanisms in developmental systems and help make Turing mechanisms more relevant to quantitative biology.
While we focus on the sweet spot of Turing network size for maximal robustness, a recent study of very large networks supports our findings, that complex random networks with $N>\!\!>1$ do not tend to spontaneously form Turing patterns [39]. When not constraining $\sigma^{2}$, this is because eigenvalue circles cross over to the positive real axis, leading to instability even without diffusion. Even when constraining the radius to $\gamma=1$, this is still the case, as we found, since the eigenvalue distribution becomes a step function, eliminating the important marginal stable cases. Remedies for overcoming above mentioned issues are identified in [39], such as allowing for non-diffusive interactions including cross-diffusion, nonlocal interactions, or advection (chemotaxis) [40].
@@
To make our model interpretable, we used a number of simplifying assumptions, resulting in limitations of our predictions. A major simplification is our diagonal matrix elements, which are set to the same value, representing equal degradation rates for all species. This led to the convenient behavior that all eigenvalues are centered around a single point on the negative real axis, following Robert May’s circular law. Without that assumption, the eigenvalue distributions would still be restricted by circles, but now follow the less restrictive Gerschgorin theorem. Based on the latter, the radius scales as $\sim N$, and not $\sim\sqrt{N}$ as in May’s circular law. This leads not only to larger circles but also to unions of overlapping circles, restricting the real parts less efficiently [41]. Using the same degradation rates also does not allow for flexibility in the levels of self-activation and inhibition often observed in biological pathways. Choosing degradation rates from a distribution introduces correlations among corresponding off-diagonal matrix elements $J_{ij}$ and $J_{ji}$ and hence different shapes of eigenvalue distributions [30].
Due to the optimal Turing network size for relatively small $N$, we did not need to investigate very large matrices. However, to speed up the investigation of large random Jacobians, the concept of reactivity can be included [42]. Reactivity, describing the transient growth potential of local perturbations in absence of diffusion, could provide a preliminary filter that avoids the full diagonalization of the system’s Jacobian. This is because the reactivity is determined by the largest eigenvalue of the Hermitian part of the Jacobian matrix, which is computationally cheaper to compute than the eigenvalues of the full Jacobian matrix.
@@
Another interesting future direction is to study random Turing networks on graphs, moving from continuous to discrete topologies, with diffusion replaced by hopping along the graph’s edges [24] [25] [26] [27]. Due to more versatile interactions, robustness may be increased. While classical Turing patterns are associated with regular spatial structures, on graphs, the lack of a physical notion of space makes patterns inherently dependent on the network’s topology. Patterns may appear irregular on random graphs or when nodes are visualized arbitrarily, but they reflect the underlying eigenmodes of the graph’s Laplacian and the topology’s influence on reaction-diffusion dynamics. The challenge lies in interpreting these patterns in the context of the system being modeled, as their biological or functional meaning may not rely on spatial regularity but rather on network-specific features. However, applications are vast, ranging from biochemical reactors and neuronal networks to tissues of living cells, synthetic systems, and patchy predator–prey ecosystems.
@@ Methods
Parameter sampling was done with Latin hypercube sampling, which is a statistical method for generating a sample of parameter values from a multidimensional space, ensuring that each parameter is evenly sampled across its range by dividing the space into equal probability intervals [43] [18]. Its strength lies in efficiently covering the input space with fewer samples compared to simple random sampling, reducing variance while maintaining representative coverage. Specifically, for Fig. 3 we sampled $10^{7}$ parameter combinations for each network topology using the lhs function from pyDOE library, using a loguniform distribution with the following bounds: $V_{i}:\ (0.1,10)$, $K_{ij}:\ (0.01,1)$, $b_{i}:\ (0.001,0.1)$, $\mu_{i}:\ (0.01,1)$ (see Eq. 2). The diffusion constants for Species 1 and 2 were fixed to $1$ and $10$, and the Hill coefficient was set to $n_{ij}=2$ throughout. For each parameter set we calculated the unique steady states using $10$ different initial guess using the Newton-Raphson method (function root with ’hybrid’ setting). In Figs. 6 and 7 we used $10^{3}$ random matrices for each $N$ and $\gamma$ pair, and in Fig. 8 we used $10^{3}$ random matrices for each $D_{1}$ and $D_{2}$ pair.
@@
Fitting a beta distribution to the data of the Jacobian matrix elements in Fig. 3 involves estimating the distribution’s parameters, typically using the method of moments or maximum likelihood estimation (MLE). The method of moments involves calculating the sample mean and variance to derive initial estimates for the shape parameters $\alpha$ and $\beta$. MLE, on the other hand, maximizes the log-likelihood function, which measures how well the beta distribution with specific parameters explains the observed data, to find the best-fitting parameters. Both methods aim to provide a beta distribution that accurately represents the underlying data distribution. Here, we used the method of moments to obtain an initial guess for the parameter values of $\alpha$ and $\beta$, which we then used as an initial guess for the maximization of the log-likelihood function.
@@
Hence, the last of Turing’s conditions cannot be fulfilled for two nodes, which is relevant for Figs. 7A and 8A. Note, Jacobian matrix Eq. LABEL:eq:J2D is different from Turing’s original linear model, in which the diagonal matrix element of the activator is positive without diffusion [4]. In our case, all diagonal matrix elements are $-1$, leading to increased stability.